OpenStudy (anonymous):

What is the force of gravitational attraction between a 20,000 kg comet and a 7000 kg asteroid which are 1000 m apart? 9.338 x 10-9 N 140,000 N 140 N

OpenStudy (shane_b):

Do you know the law of universal gravitation?

OpenStudy (anonymous):

yeah F=Gm1 m2/d2

OpenStudy (shane_b):

Yes, it's:\[\Large F_g=G\frac{m_1m_2}{d^2}\]All you need to do is plug in the values you were given.

OpenStudy (anonymous):

140000?

OpenStudy (shane_b):

Not even close :) Use a calculator

OpenStudy (shane_b):

\[\Large F_g=(6.674^{-11}Nm^2kg^{-2}(\frac{(20000kg)(7000kg)}{(1000m)^2}=?\]

OpenStudy (anonymous):

9.338 x 10-9 N

OpenStudy (shane_b):

Yes :)